C= 450 +0.4Y
I=350
G=150
X=70
Z= 35+0.1Y
T= 0.15Y
Yf= 1550

Calculate what the new equilibrium income should be if the government of this country decides to cancel all taxes, implying the tax rate would now be 0%
1
Expert's answer
2020-11-30T16:34:21-0500

"Yf=C-T+I+G+X-Z"

"1550=450+0.4Y-0.15Y+350+150+70-35-0.1Y"

"Y=3766.67" - equilibrium income with tax

"1550=450+0.4Y+350+150+70-35-0.1Y"

"Y=1883.33" - new equilibrium income without tax